OpenStudy (anonymous):

if f(x)=3lnx+C and f(e)=2 then find the value of C help please<3

OpenStudy (lgbasallote):

\[f(x) = 3\ln x + C\] \[f(e) = 2\] just substitute the x in f(x) with e and equate it to 2 \[f(e) = 3\ln (e) + C \] \[3 \ln (e) + C = 2\] \[3(1) + C = 2\] \[3 + C = 2\] \[C = 2- 3\] \[C = -1\] got it?

OpenStudy (anonymous):

oh ok but may i ask why you made lnx = 1?

OpenStudy (anonymous):

or ln(e)

OpenStudy (lgbasallote):

because ln is \(\large \log_e\) therefore \[\large \ln e = \log_e e = 1\]

OpenStudy (anonymous):

ohh ok thankss

OpenStudy (lgbasallote):

welcome ^_^